OpenStudy (anonymous):

Father' age was 7 times his son before five years, after five years he will be 3 times his son, what is the age of father?

OpenStudy (anonymous):

present age of son = X son age before 5 years =X-5 son age after 5 years = X+5

OpenStudy (anonymous):

before five years father age=7(x-5) father age at present = 7(X-5)+5 Age of father after 5 years = 3(X+5) So, present age of father = 3(X+5)-5

OpenStudy (anonymous):

after solving both present age we get x=10 that means son age is 10years and father age is 40years

OpenStudy (jack1):

thumbs up dude, perfect

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Thank you @HARSH123...